# SPIRIT: Short-term Prediction of solar IRradIance for zero-shot Transfer learning using Foundation Models

### v1

@@ 3.3. Forecasting Architecture
where $\oplus$ denotes concatenation, providing a comprehensive characterization of each sample $(X_{t},\mathbf{A}_{t},y_{t})\in\mathcal{D}$.
Thus, the collection of feature vectors over the sequence of $T$ time steps is given by:
$\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$
where $\mathbf{F_{1:T}}$ represents the set of concatenated feature representations created for each timestamp in the sequence.
Given the collection of feature vectors $\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$, a time-series model $\mathcal{M}$ is used to encode the observed sequence into a latent vector $\mathbf{L}\in\mathbb{R}^{m}$, which captures the full context of the input data series while retaining its temporal patterns and dependencies:
@@
where $\mathcal{M}$ represents the time-series model that transforms the observed sequence of feature vectors $\mathbf{F_{1:T}}$ into a compact representation in the latent space $\mathbb{R}^{m}$.
To integrate known future information, derived from the spatiotemporal context of time and location, future covariate vectors $\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$ are constructed for each forecast time $T+\tau_{i}$, . The full covariate vector $\mathbf{C}\in\mathbb{R}^{q\cdot H}$ is then formed by concatenating these individual representations across all forecast horizons:
$\mathbf{C}=\bigoplus_{i=1}^{H}\mathbf{C}_{T+\tau_{i}},\quad\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$
We concatenate the future covariate vector $\mathbf{C}$ with the latent representation of the past time steps $\mathbf{L}$, forming the final vector that encompasses all relevant information:
$\mathbf{h}=\mathbf{L}\oplus\mathbf{C}$
This ensures that both past contextual information as well as known future data contribute to the forecasting process.
Next, a regression function $R_{\omega}:\mathbb{R}^{m+q\cdot H}\to\mathbb{R}^{H}$, parameterized by $\omega$, is applied to the vector $\mathbf{h}\in\mathbb{R}^{m+q\cdot H}$ to generate the corresponding predicted GHI values. The regressor outputs a vector $\hat{\mathbf{y}}\in\mathbb{R}^{H}$ of predicted GHI values for the forecast time intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$:
@@
Forecasting loss function $\mathcal{L}_{forecast}(\omega)$ is defined as the mean of the individual regression losses computed over all forecast intervals $T+\tau_{j}$ for each sample $i$ Specifically, the total loss is given by:
@@
where $\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$ is the individual regression loss for the forecast interval $T+\tau_{j}$ for sample $i$ To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{forecast}(\omega)$ using gradient-based optimization methods. The complete architecture is illustrated in Figure 1.
@@
The Significance of Generalized Encoders: A key distinction of our approach is that in prior work (Gao and Liu, 2022; Hasan, 2023; Siddiqui et al., 2019), the encoder $E$ is a vision model typically trained on data from a specific location and camera setup. Furthermore, studies aiming for generalizability typically rely on training models using a fusion of solar datasets from multiple locations (Nie et al., 2024; Despotovic et al., 2024). In contrast, we argue, and later demonstrate, that leveraging a foundation model, a highly generalizable feature extractor, provides a more robust $E$ function. A foundation model not only matches the performance of site-specific encoders at a given location with a particular setup but also demonstrates an unparalleled advantage in generalizing across diverse locations and camera setups.
3.4. Physics-inspired Feature Engineering
Clear sky models (Ineichen and Perez, 2002; Stein et al., 2012; Perez et al., 2002; Mueller et al., 2004) are mathematical models that estimate the theoretical solar irradiance at a given location under cloud-free conditions, serving as a representation of the maximum possible radiation reaching the Earth’s surface. These models leverage fundamental atmospheric physics and employ mathematical formulations based on solar geometry (Stein et al., 2012), atmospheric transmittance (Stein et al., 2012), and radiative transfer (Stein et al., 2012) to derive estimations of GHI, DNI and DHI under clear sky conditions. The Ineichen clear sky model (Ineichen and Perez, 2002) requires inputs such as latitude, longitude, time, and date, which are readily available. This allows clear sky irradiance values to be readily computed and incorporated into our model as features, providing a reference for expected irradiance levels in the absence of cloud interference.
Physics behind solar irradiance: Solar irradiance, the power per unit area received from the Sun in the form of electromagnetic radiation, is measured in watts per square meter $(W/m^{2})$. The amount of solar irradiance received by a solar panel depends on additional site-specific factors, including the panel’s tilt and orientation angle, the Sun’s altitude and azimuth, and the geographic location’s latitude and longitude. We first look at the angle of incidence ($\theta$(Laboratories, [n. d.]), i.e. the angle between the incoming solar rays and the normal to the surface of the solar panel. It can be calculated using the following formula:
(2) $cos(\theta)=cos(\theta_{z})\cdot cos(\beta)+sin(\theta_{z})\cdot sin(\beta)\cdot cos(\gamma-\alpha)$
where $\theta_{z}$ and $\gamma$ are the solar zenith and azimuth angles respectively. While $\beta$ and $\alpha$ are the tilt and azimuth angles of the panel.
We calculate the effective irradiance by adding the three main components: direct, diffuse, and reflected irradiance (see below):
(3) $I_{panel}=DNI\cdot cos(\theta)+DHI\cdot\frac{1+cos(\beta)}{2}+GHI\cdot\rho\cdot\frac{1-cos(\beta)}{2}$
where $I_{panel}$ is the effective irradiance and $\rho$ is the ground reflectance.
4. SPIRIT Implementation
4.1. Nowcasting
In our approach, we utilize the pre-trained Google Vision Transformer (ViT) (Dosovitskiy, 2020), a model with 632 million parameters, to generate embeddings for sky camera images. To reduce sensor dependence and focus on image features, we exclude meteorological sensor data, incorporating only auxiliary variables such as zenith and azimuth angles, clear sky irradiance, panel tilt, and orientation. These image embeddings are subsequently concatenated with the auxiliary vector to form the final feature representation. The combined feature vectors, paired with their corresponding ground truth GHI values, are then used to train an XGBoost regressor within a supervised learning framework. The model is optimized by minimizing the Mean Squared Error (MSE) loss function, which measures the difference between the predicted and actual GHI values.
4.2. Forecasting
For forecasting, we employ the Google Vision Transformer (ViT) (Dosovitskiy, 2020) to generate image embeddings, which are subsequently concatenated with the auxiliary variables to form a comprehensive feature representation. To account for temporal dependencies, we input a sequence of six images, representing a 1-hour context window, into a transformer-based time-series encoder (Vaswani, 2017). This encoder processes the temporal sequence and learns a latent representation of the past context, which is then fused with a future covariate vector that includes azimuth and zenith angles, as well as clear sky GHI. The resulting representation is passed through a multi-layer perceptron (MLP) to predict the solar irradiance for the 1-hour, 2-hour, 3-hour, and 4-hour forecast intervals. This implementation exemplifies one approach in our framework, with additional variations incorporating different vision encoders of varied sizes in the ablation studies detailed in Section 7.
5. Evaluation Methodology
5.1. Datasets
@@
We evaluate our methods using three publicly available datasets: TSI880 (Andreas and Stoffel, 1981), ASI16 (Andreas and Stoffel, 1981), and SKIPP’D (Nie et al., 2023). The TSI880 and ASI16 datasets, both collected from the NREL Solar Radiation Research Laboratory in Golden, Colorado, provide sky images captured every 10 minutes along with corresponding GHI values and auxiliary data such as air temperature and relative humidity and only differ in camera setup and sensors, with the ASI16 dataset capturing higher-resolution images. The SKIPP’D dataset, collected from Stanford University, consists of raw sky images captured every minute and PV power output data, prioritizing finer temporal granularity at the expense of image quality. For more details, refer to Appendix A.
@@
We utilize the TSI880 and ASI16 datasets to investigate the impact of camera setup at the same location. To explore location and task shifts, we use the SKIPP’D dataset to evaluate the performance of models trained on GHI data in predicting PV power output. The SKIPP’D dataset features lower-resolution images and lacks meteorological data, thereby presenting a more challenging task by limiting the contextual information typically leveraged by prior models (Gao and Liu, 2022; Siddiqui et al., 2019). To ensure the models learn from higher-quality, information-rich datasets, we train exclusively on the TSI and ASI datasets while evaluation is done across all the datasets, including the more challenging SKIPP’D, allowing us to assess how well the models generalize to lower-quality data and increased domain shifts.
5.2. Performance Metrics
@@
We assess the effectiveness of the predicted values using the normalized Mean Absolute Percentage error (nMAP), defined as:
(4) $\text{nMAP}=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\frac{1}{N}\sum_{i=1}^{N}y_{i}}\times 100$
@@
where $y_{i}$ represents the actual value and $\hat{y}_{i}$ represents the predicted value for the $i$-th sample, with $i\in\{1,\dots,N\}$. It is commonly used for solar irradiance prediction as the normalization ensures that models can be assessed uniformly on datasets with varied value ranges, avoiding biased assessments due to scale differences.
5.3. Baselines
To benchmark our proposed method, we compare its performance against the state-of-the-art baseline, Gao et al. (Gao and Liu, 2022), who demonstrate state-of-the-art performance for nowcasting and forecasting by training a vision transformer (Dosovitskiy, 2020) from the ground up using 10 years of site-specific data (Andreas and Stoffel, 1981; Gao and Liu, 2022; Siddiqui et al., 2019). Their approach for forecasting utilizes a temporal transformer (Vaswani, 2017), also trained on the same duration of data. To ensure a fair comparison, we reproduced their architecture and conducted experiments under the same conditions for both Gao et al.’s (Gao and Liu, 2022) model and SPIRIT.
Table 1. Nowcasting performance across multiple datasets: SPIRIT and Gao et al.’s (Gao and Liu, 2022) model trained on one dataset for a year are evaluated with nMAP both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017. We observe comparable performance when tested in the training setup, but our model demonstrates significantly better zero-shot performance in a new location.
Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@ 5.3. Baselines
Table 2. Forecasting performance across multiple datasets and forecast intervals: SPIRIT and Gao et al.’s (Gao and Liu, 2022) model trained on one dataset are evaluated with nMAP error both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017 across four forecast intervals: 1hr, 2hr, 3hr, and 4hr.
Interval Trained on Tested on SPIRIT Gao et al. (Gao and Liu, 2022)
@@
5.4. Zero-shot Transfer Learning
@@
To evaluate the zero-shot generalization performance of our models, we analyze two distinct transfer learning scenarios. The first scenario examines intra-location generalization, where the models are trained and tested in the same geographic location but under varying camera setups. While the environmental conditions remain consistent, variations in camera setup, viewing angles, and image resolutions exist between the training and testing phases. When image-based models are trained on data from a particular camera setup, they learn to associate specific regions of the image with key features—such as the position of the sun, cloud formations, or atmospheric conditions—that influence the predicted output. However, when the camera setup is altered, the spatial mapping of these features within the image shifts. To assess how well the models handle such variations, we train them using the TSI dataset and evaluate them on the ASI dataset, and vice versa.
@@
The second scenario focuses on cross-location and cross-task generalization, where models trained in one geographic location are tested in another with different environmental and sensor characteristics. We train on the TSI and ASI datasets and evaluate on the SKIPP’D dataset, with the task shifting from predicting GHI to PV power output. Since GHI and PV output have a nearly linear correlation (Vilanova et al., 2020), this serves as a valid example of heterogeneous transfer learning. To account for the significant scale difference between GHI and PV output, model outputs are normalized for comparability. We conduct experiments for both nowcasting and forecasting tasks, training the models for one year and testing them on another year to account for seasonal variations, thus ensuring a fair evaluation. The nMAP errors are reported in Table 1 for nowcasting and Table 2 for forecasting, comparing SPIRIT with the state-of-the-art in both-the zero-shot transfer learning setups and the traditional settings, where the models are trained and tested using data from the same location and setup, but different years.
Figure 2. We compare the nowcasting performance of SPIRIT and Gao et al. using nMAP error. The solid lines represent the average performance across different fine-tuning training sizes, measured in weeks of data. The shaded regions indicate the xconfidence interval, reflecting variability across multiple experimental settings, including training on one dataset and testing on another, as well as selecting fine-tuning data from different starting points throughout the year.
@@
Figure 3. We compare the forecasting performance of SPIRIT and Gao et al. using nMAP error across different forecast intervals. Subfigures (a), (b), (c), and (d) correspond to 1-hour, 2-hour, 3-hour, and 4-hour forecasting, respectively. The solid lines represent the average performance for each forecast interval, with varying fine-tuning training sizes measured in weeks of data. The shaded regions denote the 95% confidence interval, illustrating the variability across multiple experimental settings, including training on one dataset and testing on another, as well as selecting weeks of contiguous fine-tuning data from different starting points throughout the year. SPIRIT exhibits consistently low variance compared to the baseline, particularly in settings with severely limited data, demonstrating its ability to maintain stability. In contrast, the baseline shows high variance, indicating uncertainty in its predictions.
5.5. Fine-tuning with Limited Data
Building upon our zero-shot transfer learning experiments, we now investigate the adaptability of our models in a fine-tuning framework, where a limited amount of labeled data from the target domain is available for fine-tuning. This scenario closely resembles practical deployment conditions, where prolonged data collection is often infeasible, and models must quickly adapt to new locations with minimal supervision. We evaluate transfer learning with limited data in two scenarios: intra-location adaptation and cross-location adaptation, as in Subsection 5.4.
For both experimental setups, we perform fine-tuning using progressively increasing amounts of labeled data from the target domain—specifically, one, two, three, and four weeks of data from a full year for nowcasting, and two, four, eight, twelve, and sixteen weeks of data for forecasting with testing done on the remaining data from the year. Given the greater complexity of forecasting, we extend the fine-tuning experiment to a larger time frame. Additionally, due to the requirement for temporal consistency in the time series data, as discussed in Subsection A.2, the number of nowcasting samples for a given time period exceeds that of forecasting samples. We implement a selective fine-tuning approach, where only the regressors (see Figure 1) are updated, while the rest of the model is frozen. This ensures that the pre-trained feature representations, which capture generalizable spatiotemporal patterns, remain intact while allowing the model to adapt to location- and camera-specific variations. As demonstrated in prior work (Nie et al., 2024; Zhou et al., 2020; Sarmas et al., 2022), fine-tuning only the final layers achieves competitive adaptation performance while mitigating the risk of overfitting to the limited target data The nowcasting metrics are shown in Figure 2, and the forecasting metrics are depicted in Figure 3.
6. Results
6.1. Zero-shot Transfer Learning

### v2

@@ 3.4 Forecasting Architecture
where $\oplus$ denotes concatenation, providing a comprehensive characterization of each sample $(X_{t},\mathbf{A}_{t},y_{t})\in\mathcal{D}$. Thus, the set of feature vectors for each timestamp over the sequence of $T$ time steps is given by $\mathbf{F_{1:T}}=\{\mathbf{f}_{1},\mathbf{f}_{2},\dots,\mathbf{f}_{T}\}$.
A time-series model $\mathcal{M}$ is used to encode the observed sequence $\mathbf{F_{1:T}}$ into a latent vector $\mathbf{L}\in\mathbb{R}^{m}$, which captures the full context of the input data series while retaining its temporal patterns and dependencies:
@@
To integrate known future information, derived from the spatiotemporal context of time and location, future covariate vectors $\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$ are constructed for each forecast time $T+\tau_{i}$. The full covariate vector $\mathbf{C}\in\mathbb{R}^{q\cdot H}$ is then formed by concatenating these individual representations across all $H$ forecast horizons:
$\mathbf{C}=\mathop{\mathchoice{\displaystyle\oplus}{\textstyle\oplus}{\scriptstyle\oplus}{\scriptscriptstyle\oplus}}_{i=1}^{H}\mathbf{C}_{T+\tau_{i}},\quad\mathbf{C}_{T+\tau_{i}}\in\mathbb{R}^{q}$
We concatenate the future covariate vector $\mathbf{C}$ with the latent representation of the past time steps $\mathbf{L}$, forming the final vector $\mathbf{h}$ that encompasses all relevant information: $\mathbf{h}=\mathbf{L}\oplus\mathbf{C}$. This ensures that both past contextual information as well as known future data contribute to the forecasting process. A regression function $R_{\omega}:\mathbb{R}^{m+q\cdot H}\to\mathbb{R}^{H}$, parameterized by $\omega$, is applied to the vector $\mathbf{h}\in\mathbb{R}^{m+q\cdot H}$ to generate the corresponding predicted GHI values. The regressor outputs a vector $\hat{\mathbf{y}}\in\mathbb{R}^{H}$ of predicted GHI values for the forecast time intervals $T+\tau_{1},T+\tau_{2},\dots,T+\tau_{H}$:
@@
Forecasting loss function $\mathcal{L}_{forecast}(\omega)$ is defined as the mean of the individual regression losses computed over all forecast intervals $T+\tau_{j}$ for each sample $i$ across $N$ samples and $H$ forecast horizons. The total loss is given by:
@@
where $\mathcal{L}(\hat{y}^{(i)}_{T+\tau_{j}},y^{(i)}_{T+\tau_{j}})$ is the individual regression loss for the forecast interval $T+\tau_{j}$ for sample $i$ . To learn the optimal parameters $\omega^{*}$, we minimize $\mathcal{L}_{forecast}(\omega)$ using gradient-based optimization methods. The complete architecture is illustrated in Figure 1.
@@
The Significance of Generalized Encoders: A key distinction of our approach is that in prior work (wacv2022; prior_methods_site_specific; talha2019), the encoder $E$ is a vision model typically trained on data from a specific location and camera setup. Furthermore, studies aiming for generalizability typically rely on training models using a fusion of solar datasets from multiple locations (yuhao_transfer_learning; site_specific_trained_transfer_learning). In contrast, we argue, and later demonstrate, that leveraging a foundation model, a highly generalizable feature extractor, provides a more robust $E$ function. A foundation model not only matches the performance of site-specific encoders at a given location with a particular setup but also demonstrates an unparalleled advantage in generalizing across diverse locations and camera setups.
3.5 Implementation
Our implementation leverages the pre-trained Google ViT (google_vit) for image embedding generation, deliberately excluding all meteorological sensor data to significantly reduce sensor dependence. Instead, we incorporate only the zenith and azimuth angles and, physics-based features (clear sky irradiance values, panel tilt and orientation). These carefully selected image embeddings are concatenated with the auxiliary vector to form robust feature representations that train an XGBoost regressor for nowcasting. For forecasting, we extend these feature representations across a one-hour context window (six consecutive timestamps) and process them through a transformer encoder (vaswani2017attention). The resulting temporal representation is fused with future covariates from the auxiliary and physics features before passing through MLP layers to predict solar irradiance at 1-4 hour horizons, using clear-sky GHI as residuals before the final regressor. This represents one concrete realization of our system, with comprehensive ablations in Appendix A for understanding the contributions of each component. More details on hyperparameters and deployment efficiency are provided in Appendix F.
4 Evaluation Methodology
4.1 Datasets and Performance Metrics
@@
We evaluate our methods using three publicly available datasets: TSI880 (tsi_dataset), ASI16 (tsi_dataset), and SKIPP’D (SKIPPD_dataset). The TSI880 and ASI16 datasets, both collected from the NREL Solar Radiation Research Laboratory in Golden, Colorado, provide sky images captured every 10 minutes along with corresponding GHI values and auxiliary data such as air temperature and relative humidity and only differ in camera setup and sensors, with the ASI16 dataset capturing higher-resolution images. The SKIPP’D dataset, collected from Stanford University, consists of raw sky images captured every minute and PV power output data, prioritizing finer temporal granularity at the expense of image quality. For more details, refer to Appendix B.
@@
We utilize the TSI880 and ASI16 datasets to investigate the impact of camera setup at the same location. To explore location and task shifts, we use the SKIPP’D dataset to evaluate the performance of models trained on GHI data in predicting PV power output. The SKIPP’D dataset features lower-resolution images and lacks meteorological data, thereby presenting a more challenging task by limiting the contextual information typically leveraged by prior models (wacv2022; talha2019). To ensure the models learn from higher-quality, information-rich datasets, we train exclusively on the TSI and ASI datasets while evaluation is done across all the datasets, including the more challenging SKIPP’D, allowing us to assess how well the models generalize to lower-quality data and increased domain shifts.
@@
We assess the effectiveness of the predicted values using the normalized Mean Absolute Percentage error (nMAP):
$\text{nMAP}=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\frac{1}{N}\sum_{i=1}^{N}y_{i}}\times 100$ (2)
@@
where $y_{i}$ represents the actual value and $\hat{y}_{i}$ represents the predicted value for the $i$-th sample, with $i\in\{1,\dots,N\}$. It is commonly used for solar irradiance prediction as the normalization ensures that models can be assessed uniformly on different datasets, avoiding biased assessments due to scale differences. Comprehensive analysis on metrics is in Appendix I.
4.2 Baselines
We benchmark against wacv2022, the current state-of-the-art, who achieve superior performance by training a vision transformer (google_vit) and temporal transformer (vaswani2017attention) on 10 years of site-specific data (tsi_dataset; wacv2022; talha2019) with sensor-dependent auxiliary values. For equitable evaluation, we reproduced their architecture and conducted experiments under identical conditions for both their model and SPIRIT. More details on baselines are in Appendix G.
Train Test SPIRIT wacv2022
@@ 4.2 Baselines
Table 1: Nowcasting performance across multiple datasets: SPIRIT and wacv2022 model trained on one dataset for a year are evaluated with nMAP error (lower is better) both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017. Performance is comparable when tested in the training setup, but our model demonstrates significantly better zero-shot performance when in a new setup or a new location.
Interval Train Test SPIRIT wacv2022
@@
Table 2: Forecasting performance across multiple datasets and forecast intervals: SPIRIT and wacv2022 model trained on one dataset are evaluated with nMAP error both in a zero-shot setting and on the same dataset, with testing on TSI 2021, ASI 2021, and SKIPP’D 2017 across four forecast intervals: 1hr, 2hr, 3hr, and 4hr.
4.3 Zero-shot Transfer Learning
@@
To evaluate the zero-shot generalization performance of our models, we analyze two distinct transfer learning scenarios. First, we examine intra-location generalization where models are trained and tested at identical geographic locations but with differing camera configurations. Though environmental conditions remain consistent, variations in camera setup alter the spatial mapping of critical features within images. Models trained on specific camera setups learn to associate image regions with key elements (sun position, cloud formations, atmospheric conditions) that influence predictions. When these spatial relationships shift due to altered camera configurations, model performance deteriorates. We evaluate this scenario by training on TSI and testing on ASI (both from Golden, Colorado), and vice versa.
@@
Second, we investigate cross-location and cross-task generalization, training models in one geographic location and testing in another with different environmental and sensor characteristics. Specifically, we train on TSI and ASI datasets and evaluate on SKIPP’D (from Stanford University, California), shifting from GHI prediction to PV power output forecasting. Given the nearly linear correlation between GHI and PV output (ghi_pv_linear), this constitutes a valid heterogeneous transfer learning scenario. Model outputs are normalized to address scale differences between GHI and PV measurements, as detailed in Appendix F.2. For comprehensive evaluation, we conduct experiments for both nowcasting and forecasting tasks, training on one year and testing on another to account for seasonal variations. Tables 1 and 2 report nMAP errors for nowcasting and forecasting respectively, comparing SPIRIT against the state-of-the-art in both zero-shot transfer learning and traditional setting (same location and setup).
Figure 2: Mean nMAP performance of SPIRIT and wacv2022 across varying fine-tuning data sizes in weeks. Solid lines denote average accuracy and shaded regions show 95% confidence intervals across multiple runs, including cross-setup (TSI to ASI and vice versa) and cross-location (TSI or ASI to SKIPP’D) transfer scenarios.
@@
Figure 3: Forecasting performance of SPIRIT and wacv2022 using nMAP error across different forecast intervals. Subfigures (a), (b), (c), and (d) correspond to 1-hour, 2-hour, 3-hour, and 4-hour forecasting, respectively. The solid lines represent the average performance, with varying fine-tuning data sizes (in weeks). The shaded regions denote the 95% confidence interval, illustrating the variability across multiple runs, including training on one dataset and finetuning and testing on another, as well as utilizing randomized seasonal sampling. SPIRIT exhibits consistently better performance and low variance compared to the baseline, particularly with severely limited data, demonstrating its ability to maintain stability over the baseline’s.
4.4 Fine-tuning with Limited Data
Following our zero-shot transfer learning experiments, we examine SPIRIT’s adaptability through fine-tuning with limited target domain data. This scenario replicates real-world deployment constraints where extensive data collection is impractical and models must adapt rapidly with minimal supervision. We evaluate both intra-location and cross-location/cross-task adaptation scenarios as described in Subsection 4.3. For both experimental configurations, we incrementally increase the amount of target domain data used for fine-tuning, one to four weeks for nowcasting and two to sixteen weeks for forecasting, with testing performance on the remaining annual data. The longer timeframe for forecasting accounts for the reduced sample density due to temporal consistency constraints, highlighted in Appendix B.3, and forecasting’s fundamentally higher complexity relative to nowcasting.
We employ a selective fine-tuning strategy wherein only the regressor layers (Figure 1) are updated while freezing the remainder of the temporal encoder. This preserves the pre-trained feature extractors that capture generalizable spatiotemporal patterns, an essential part of our core hypothesis. Selective fine-tuning of terminal layers not only accelerates adaptation but also mitigates overfitting on limited target data (yuhao_transfer_learning; uses_freezing_for_tl; transfer_learning_strategies_comparison). Detailed rationale for this choice is provided in Appendix D.3. The nowcasting and forecasting performance metrics are presented in Figures 2 and 3, respectively.
5 Results
5.1 Zero-shot Transfer Learning